A car’s speed can be calculated by the equation d=st, where d is distance (meters), s is speed in (m/s), and t is time (sec). What is the speed of the car if the distance traveled is 30km in 20 minutes?
Answer #1

We know that:

d = s*t

distance = speed*time

speed = distance/time

distance traveled by car = d = 30 km = 30000 m

Since 1 km = 1000 m

time taken = 20 min = 1200 sec

Since 1 min = 60 sec

So,

Speed = 30000 m/1200 s

speed = (30000/1200) m/s = (300/12) m/s

Speed = 25 m/s

Let me know if you've any query.